# Memory Mosaics

### v2

@@ Abstract
Memory Mosaics are networks of associative memories working in concert to achieve a prediction task of interest. Like transformers, memory mosaics possess compositional capabilities and in-context learning capabilities. Unlike transformers, memory mosaics achieve these capabilities in comparatively transparent ways. We demonstrate these capabilities on toy examples and we also show that memory mosaics perform as well or better than transformers on medium-scale language modeling tasks.
[…]
@@ 1 Introduction
This paper presents a learning system architecture, Memory Mosaics, in which multiple associative memories work in concert to carry out a prediction task of interest. Such systems are closely related to memory networks (Weston et al., 2014; Sukhbaatar et al., 2015) and resemble transformers (Vaswani et al., 2017) despite significant differences. Like transformers, Memory Mosaics possesses some of the disentanglement and compositional capabilities that have long eluded machine learning systems (Lake and Baroni, 2018). Unlike transformers whose internal mechanism are hard to decipher (Olsson et al., 2022; Bietti et al., 2024), Memory Mosaics achieve these capabilities in comparatively transparent ways.
The three main contributions of this work are (a) recognizing and exploiting a similarity between smoothing associative memories and self-attention, (b) identifying and illustrating the predictive disentanglement principle which explains how training decomposes the overall task in interesting ways, and (c) showing that this comparatively transparent architecture matches the performance of decoding transformers on a language modeling task.
Section 2 describes the basic architecture and outlines its consequences. Section 3 illustrates the predictive disentanglement principle. Section 4 extends these ideas to fully formed memory mosaics. Section 5 reports on medium-scale language modeling experiments.
@@ 2 Memories
$f\big{(}k;\>\{(k_{1},v_{1})\dots(k_{n},v_{n})\}\big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \mathbb{E}\/(V\>|\>K=k)\,.$ (1)
Such a conditional expectation can be constructed with Gaussian kernel smoothing, [^1] This expression results from estimating the key-value distribution with Kernel Density Estimation using a Gaussian kernel. The Parzen windows convergence theorem then ensures that (2) converges to the true conditional expectation when $n\rightarrow\infty$ and $\beta=\sqrt{n}$.
$f\big{(}k;\>\{(k_{1},v_{1})\dots(k_{n},v_{n})\}\big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{i=1}^{n}\frac{1}{Z}\leavevmode\nobreak\ e^{-\beta\|k-k_{i}\|^{2}}v_{i}\leavevmode\nobreak\ \quad\text{with}\quad Z=\sum_{i=1}^{n}e^{-\beta\|k-k_{i}\|^{2}}\,.$ (2)
The close connection between kernel smoothing and attention (Bahdanau et al., 2015) is particularly obvious when all key vectors $k_{i}$ share a same squared norm because expression (2) can be simplified as
$f\big{(}k;\>\{(k_{1},v_{1})\dots(k_{n},v_{n})\}\big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{i=1}^{n}\leavevmode\nobreak\ \frac{e^{\,\beta\,k^{\!\top}k_{i}}}{\sum_{j=1}^{n}e^{\,\beta\,k^{\!\top}k_{j}}}\leavevmode\nobreak\ v_{i}\leavevmode\nobreak\ .$ (3)
There are of course more advantageous ways to implement such an associative memory, ranging from fast Gauss transforms (Greengard and Strain, 1991; Yang et al., 2003) to locality sensitive hashing (Spring and Shrivastava, 2017; Chen et al., 2020). Although these methods will certainly prove useful in the future, this paper only relies on associative memories implemented with Gaussian kernel smoothing, not least because that makes it easy to compute gradients.
@@
Figure 1: Elementary memory unit. The keys $k_{T}$ are computed as a function of past observations $(x_{t})_{t\leq T}$. The values $v_{T}$ peek into the future. In this example, the value also depend on the next observation $x_{T+1}$. At time $T$, the associative memory uses the known key $k_{T}$ to retrieve an approximation $y_{T}$ of the yet unknown value $v_{T}$ using the previously stored pairs $(k_{t},v_{t})$. At time ${T+1}$, the value $v_{T}$ can be computed, and the pair $(k_{T},v_{T})$ can be stored into the memory.
@@
$v_{T}=\psi({\color[rgb]{.75,0,.25}x_{T+1}},x_{T},x_{T-1},\dots;{\mathbf{w}})\,.$ (5)
At any given time $T$, the associative memory contains the previously observed pairs $(k_{t},v_{t})_{t\leq{T-1}}$ and retrieves an estimate $y_{T}$ of the yet unknown value $v_{T}$ by querying the freshly computed key $k_{T}$. One time step later, the observation $x_{T+1}$ becomes available, the value $v_{T}$ can be computed, and the pair $(k_{T},v_{T})$ is stored into the memory.
The exact form of the feature extraction functions can vary considerably in complexity. For instance, when each observation $x_{T}$ carries sufficient information, the keys $k_{T}$ and values $v_{T}$ can be computed as linear functions of respectively $x_{T}$ and $x_{T+1}$
$k_{T}$ $=W_{\varphi}\,x_{T}$ (6)
$v_{T}$ $=W_{\psi}\,x_{T+1}\leavevmode\nobreak\ .$
However it is often attractive to consider more complicated feature extraction functions involving convolutions, leaky averages, layers, transfer functions, normalization, or even polynomial time-series summaries (Voelker et al., 2019; Gu et al., 2020). For instance, the language experiments (Section 5) use feature extractors of the following form, with $\mathrm{Norm}(x)=x/\|x\|$:
$k_{T}$ $=\alpha_{\varphi}\mathrm{Norm}\big{(}{\color[rgb]{.75,0,.25}\bar{k}_{T}}\big{)}$ $\text{with}\quad\overbrace{{\color[rgb]{.75,0,.25}\bar{k}_{T}}={\color[rgb]{0,0,1}\tilde{k}_{T}}+\lambda_{\varphi}{\color[rgb]{.75,0,.25}\bar{k}_{T-1}}\quad\quad{\color[rgb]{0,0,1}\tilde{k}_{T}}=W_{\varphi}\,x_{T}}^{\text{leaky average over t = T, {T-1}\ldots, 1}}$ (7)
$v_{T}$ $=\alpha_{\psi}\mathrm{Norm}\big{(}{\color[rgb]{.75,0,.25}\bar{v}_{T}}\big{)}$ $\text{with}\quad\underbrace{{\color[rgb]{.75,0,.25}\bar{v}_{T}}={\color[rgb]{0,0,1}\tilde{v}_{T}}+\lambda_{\psi}{\color[rgb]{0,0,1}\tilde{v}_{T+1}}\quad\quad{\color[rgb]{0,0,1}\tilde{v}_{T}}=W_{\psi}\,x_{T}}_{\text{convolution over t=T and T+1}}$
@@
Consider now a deep network that contains both glue layers and elementary memory units, with trainable weights in both the glue layers and in the feature extractors of the memory units. Together, these weights determine what each memory unit memorizes (through the key and value extraction functions) and how their outputs are combined to perform the overall prediction task.
@@
When the associative memories are implemented with differentiable kernel smoothing mechanisms, training such a deep network is simply a matter of unrolling the network in time and back-propagating the gradients, in ways that users of modern deep learning software will find very familiar. Unsurprisingly, unrolling equation (3) along an input sequence $(x_{1}\dots x_{D})$ of duration $D$ yields an expression that very much resembles masked self-attention (Vaswani et al., 2017).
$\forall\,T\in\{1\dots D\}\qquad y_{T}=\sum_{i=1}^{T-\delta}\leavevmode\nobreak\ \frac{e^{\beta\,k_{T}^{\!\top}k_{i}}}{\sum_{j=1}^{T-\delta}e^{\beta\,k_{T}^{\!\top}k_{j}}}\leavevmode\nobreak\ v_{i}\leavevmode\nobreak\ ,$ (8)
There are however notable differences between (8) and classical self-attention:
  • A more aggressive masking involves the number $\delta\geq 1$ of future time steps needed to compute $v_{T}$.
  • There is no need for position encoding of for distinction between key vectors and query vectors.
  • The eventual purpose of each memory unit is observable because, at each time $T$, its outputs $y_{T}$ can be interpreted as a conditional expectation (1) that predicts the yet unknown but explicit quantity $v_{T}=\psi\big{(}(x_{t})_{t\leq{T+\delta}};\,{\mathbf{w}}\big{)}$ on the basis of past observations.
Learning and meta-learning
This training procedure should in fact be viewed as a meta-learning process, distinct from the learning that occurs at inference time when new key-value pairs are stored into the memories. Unrolling the network in time over a complete input sequence $(x_{1}\dots x_{D})$ exposes the operation of the memories. Since the overall prediction cost averages terms evaluated at each time step along the training sequences, the training process favors statistically efficient memories, able to output useful value estimates $y_{T}$ after storing as few key-value pairs as possible (Figure 2).
@@
  • Let us first assume that each memory unit has a frozen value extraction function $\psi$. The training procedure can still make each memory unit statistically more efficient by tuning the parameters of the key extraction function $\varphi$, that is, by learning how to compare the current prediction context $(x_{T},x_{T-1},x_{T-2}\dots)$ with past prediction contexts $(x_{t},x_{t-1},x_{t-2}\dots)$ for ${t<T}$.
@@
    Learning a similarity metric —a kernel— is a well known way to make non-parametric estimators more efficient (e.g., Bach et al., 2004). For instance, the training procedure can construct keys that summarize the relevant contextual information, discarding noise factors that could increase the distance between keys associated with similar values. It can also adjust the effective kernel bandwidth, for instance, using parameter $\alpha_{\varphi}$ in equation (7).
@@
  • When multiple memory units are available, the training procedure can also optimize the value extraction functions $\psi$, redistributing the overall prediction work among the available memory units. For instance, the training procedures can construct values $v_{T}$ that are more efficiently approximated by their respective memory units, but whose approximations can still be combined to address the overall prediction task.
    We argue in this contribution that this redistribution does not only make the memory units more efficient, but also disentangles the original prediction problem in elementary sub-problems that can later be recomposed in myriad ways.
Predictive disentanglement
We have argued that training the time-unrolled network of memory units is in fact a meta-training process that learns how to efficiently use memory units to address the task of interest. Using multiple memory units rather than a single one is beneficial when the overall prediction task can be divided into smaller prediction tasks whose targets are expressed by the values $v_{T}$ and can be more efficiently predicted in isolation than together. The disentangled memories can then be recombined to provide predictions for inputs that are globally very different from the training inputs, but whose disentangled components are individually predictable. Section 3 provides a toy example.
@@
Disentanglement has long been identified as both desirable and hard to pinpoint (Bengio, 2013). Statistical definitions lack in robustness with respect to changing data distributions (Comon, 1994; Roth et al., 2022). Causal definitions cannot be tested without active experiments (Thomas et al., 2018; Bengio et al., 2019). Predictive disentanglement provides an appealing alternative that does not require a specific training algorithm but arises as a side effect of the meta-training process. Although predictive disentanglement is easier to understand in the case of a network of associative memory units, we can safely conjecture that similar phenomenon also occurs in transformers, albeit in a far less explicit manner.
@@ 3 Tracking three moons
Three moons orbit a remote planet. Although the local astronomers are very far from understanding celestial mechanics, [^2] We do not seek to discuss subtleties such as elliptical orbits or multi-body problems. Our astronomers are best compared to the ancient sky watchers whose efforts eventually gave the Ptolemaic model. they nevertheless observe periodic motions and debate how to predict future moon positions. A first astronomer proposes to compile a single table containing the daily positions of all three moons, arguing that if the current set of moon positions matches a previous observation, the future moon positions will match the following observations. A second astronomer suggests instead to make three tables, one for each moon, arguing that the future positions of each moon can be independently predicted by matching its current position with a previously observed one.
@@
$N_{h}$ $=1\text{\leavevmode\nobreak\ or\leavevmode\nobreak\ }3$
@@
$y_{t}^{(h)}$ $=\frac{1}{Z_{T}}\,\sum_{t<T}e^{\beta\,k_{T}^{(h)}\boldsymbol{\cdot}k_{t}^{(h)}}v_{t}^{(h)}$
@@
For our purposes, each observation $x_{t}$ consists of three complex numbers $e^{i\theta_{k}}$ that encode the angular positions $\theta_{k}$ of the three moons inside their respective orbital plane. We consider two single layer models (Figure 3) with either $N_{h}=1$ or $N_{h}=3$ memory units whose added dimensions match the input dimension. The key and value extraction functions follow (6) with trainable parameters collected in two ${3\times 3}$ complex matrices $W_{\varphi}$ and $W_{\psi}$. The memory unit follow (3) with a fixed parameter $\beta=50$. A third ${3\times 3}$ complex matrix $W_{z}$ combines the memory unit predictions into an output $z_{T}$ that hopefully predicts $x_{T+1}$. Both networks share an interesting analytic solution: setting all three matrices $W_{\varphi}$, $W_{\psi}$, and $W_{z}$ to the identity yields optimal predictions once the associative memories have seen enough samples.
Figure 4: Single head network prediction error versus context length. The prediction error shows a sharp transition after $\mathrm{lcm}(p_{1},p_{2},p_{3})$ observations (red vertical line), when the network switches from predicting the future moon position by repeating the last observation to predicting by find a matching memorized moon configuration.
Figure 5: Three-heads network prediction error versus context length. The prediction error improves whenever the context length becomes large enough to fully describe the orbit of each planet. The net yields accurate predictions after only $\max(p_{1},p_{2},p_{2})$ observations (last black vertical line), that is, before having seen the full set of moon configurations (red vertical line). This is possible because these unseen moon configurations are predicted by composing individual moon predictions.
@@
Figure 4 and 5 show the prediction errors of both networks as a function of the context length, that is, the number of observations stored into the memories. More precisely, for each sequence $(x_{t})$ and each time index $T$, we compute the average absolute deviation between the next 25 true moon positions $x_{T+1}\dots x_{T+25}$ and the next 25 auto-regressive predictions (in which the successive predictions are looped back into the network input.) The plots show curves averaged over 512 sequences sharing the same set of moon periods taken from either the training set or the validation set.
@@
  • For the three-heads network (Figure 5), the prediction error curve drops after seeing exactly $p_{1}$, $p_{2}$, and $p_{3}$ observations (black lines), that is whenever the orbit of an additional moon has been memorized. The network produces accurate predictions after a time equal to the period $\max(p_{1},p_{2},p_{2})$ of the slowest moon (last black line), long before the combined period $\mathrm{lcm}(p_{1},p_{2},p_{3})$ (red line) of the moon system. During this interval, accurate predictions are returned for moon configurations that can be very different from the previously observed ones. This is possible because the network combines individual moon predictions, each well supported by the past observations.
Figure 6 depicts the learned weight matrices and shows how the network has successfully disentangled the three moons ($W_{\psi}$ and $W_{\varphi}$) and learned how to recombine individual predictions ($W_{z}$).
Training notes

### v3

@@ Abstract
Memory Mosaics are networks of associative memories working in concert to achieve a prediction task of interest. Like transformers, memory mosaics possess compositional capabilities and in-context learning capabilities. Unlike transformers, memory mosaics achieve these capabilities in comparatively transparent way (“predictive disentanglement”). We illustrate these capabilities on a toy example and also show that memory mosaics perform as well or better than transformers on medium-scale language modeling tasks.
@@ 1 Introduction
This paper presents a learning system architecture, Memory Mosaics, in which multiple associative memories work in concert to carry out a prediction task of interest. Such systems are closely related to memory networks (Weston et al., 2014; Sukhbaatar et al., 2015) and resemble transformers (Vaswani et al., 2017) despite significant differences. Like transformers, Memory Mosaics possesses some of the disentanglement and compositional capabilities that have long eluded machine learning systems (Lake & Baroni, 2018). Unlike transformers whose internal mechanism are hard to decipher (Olsson et al., 2022; Bietti et al., 2024), Memory Mosaics achieve these capabilities in comparatively transparent ways.
The three main contributions of this work are (a) defining an architecture that exploits the direct similarity between self-attention and associative memories implemented with kernel regression, (b) identifying and illustrating the predictive disentanglement principle which explains how training decomposes the overall task in interesting ways, and (c) showing that this comparatively transparent architecture matches the i.i.d. performance of decoding transformers on a language modeling task, and outperforms them on o.o.d. tasks such as in-context learning.
Section 2 reviews related work. Section 3 describes simple associative memory units than can be inserted in a deep network. Section 4 explains how training such a network splits a prediction task into disentangled sub-tasks. Section 5 illustrates this “predictive disentanglement” using a network with only 54 parameters, showing that this is not a mysterious effect of scale but a property of the architecture. Section 6 extends these ideas to fully formed memory mosaics. Section 7 reports on medium-scale language modeling experiments.
2 Related Work
Several recent papers (e.g., Katharopoulos et al., 2020; Peng et al., 2023; Sun et al., 2023; Gu & Dao, 2023) propose transformer alternatives that use efficient recurrences to cut the quadratic computational cost of transformers. Closer to our interests, other authors (e.g., Ramsauer et al., 2020; Krotov, 2023; Hoover et al., 2024) rethink transformers with Hopfield-style associative memories and their associated energy function. In contrast, we leverage elementary associative memories that interpolate stored key/value pairs with a kernel regression (therefore incurring a quadratic runtime cost) in order to construct an architecture that remains very close to standard transformers but cast a new light on properties that play an important role in their compositional learning capabilities.
Closely related to predictive disentanglement, (Bengio et al., 2019) proposes a meta-learning training objective that achieves causal disentanglement by seeking quick adaptation to new distributions. We argue that a similar effect happens in our architecture, as a consequence of the normal training process interpreted as a meta-learning process, revealing an important aspect of the still mysterious compositional learning abilities of transformer-like architectures.
@@ 3 Memories
$f\big(k;\>\{(k_{1},v_{1})\dots(k_{n},v_{n})\}\big)\penalty 10000\ =\penalty 10000\ \mathbb{E}\/(V\>|\>K=k)\,.$ (1)
Such a conditional expectation can be constructed with Gaussian kernel regression, [^1] Expression (2) is known as the Nadaraya-Watson estimator (Nadaraya, 1964; Watson, 1964). It is known to converge to the true conditional expectation $\mathbb{E}(K|V)$ when $n\rightarrow\infty$ and $\beta=\sqrt{n}$.
$f\big(k;\>\{(k_{1},v_{1})\dots(k_{n},v_{n})\}\big)\penalty 10000\ =\penalty 10000\ {\sum_{i=1}^{n}\frac{1}{Z}}\penalty 10000\ e^{-\beta\|k-k_{i}\|^{2}}v_{i}\penalty 10000\ \quad\text{with}\quad Z={\sum_{i=1}^{n}}e^{-\beta\|k-k_{i}\|^{2}}\,.$ (2)
The close connection between this Gaussian kernel smoothing and attention (Bahdanau et al., 2015) is obvious when all key vectors $k_{i}$ share a same squared norm because expression (2) becomes
$f\big(k;\>\{(k_{1},v_{1})\dots(k_{n},v_{n})\}\big)\penalty 10000\ =\penalty 10000\ \sum_{i=1}^{n}\penalty 10000\ \frac{e^{\,\beta\,k^{\!\top}k_{i}}}{\sum_{j=1}^{n}e^{\,\beta\,k^{\!\top}k_{j}}}\penalty 10000\ v_{i}\penalty 10000\ .$ (3)
There are of course more advantageous ways to implement associative memories. Although some will certainly prove useful in the future, this paper only relies on associative memories implemented with Gaussian kernel smoothing, not least because that makes it easy to compute gradients.
@@
Figure 1: Elementary memory unit. The keys $k_{T}$ are computed as a function of past observations $(x_{t})_{t\leq T}$. The values $v_{T}$ peek into the future. In this example, the value also depend on the next observation $x_{T+1}$. At time $T$, the associative memory uses the known key $k_{T}$ to compute an estimate $y_{T}$ of $\mathbb{E}(v_{T}|k_{T})$ using only the previously stored pairs $(k_{t},v_{t})$, $t<T$. One time step later, the input $x_{T+1}$ is revealed, the value $v_{T}$ can be computed, and the pair $(k_{T},v_{T})$ is added to the memory.
@@
$v_{T}=\psi({\color[rgb]{.75,0,.25}\mathbf{x_{T+1}}},x_{T},x_{T-1},\dots;{\mathbf{w}})\,.$ (5)
The memory units operate independently at inference time. They start empty at the beginning of each input sequence. At time step $T$, each memory receives a key vector $k_{T}$ computed from the recent inputs $(x_{T},x_{T-1},\dots)$ and interpolates a response $y_{t}$ on the basis of the previously stored key/value pairs. The value $v_{T}$ is computed one time step later when the next input $x_{T+1}$ is revealed and the pair $(k_{T},v_{T})$ is added to the memory.
Although the value $v_{T}$ depends on the near future, the output $y_{T}$ does not depend on $v_{T}$ but merely leverages the previously stored key/value pairs to estimate $v_{T}$. Therefore there is no leak of future information: each memory unit is a little machine that predicts a bit of future information (described by $v_{T}$) on the basis of recent information (described by $k_{T}$) and previously stored key/values pairs.
The exact form of the feature extraction functions can vary in complexity. For instance, when each observation $x_{T}$ carries sufficient information, the keys $k_{T}$ and values $v_{T}$ can be computed as linear functions of respectively $x_{T}$ and $x_{T+1}$, that is $k_{T}=W_{\varphi}\,x_{T}$ and $v_{T}=W_{\psi}\,x_{T+1}$. However we find useful to consider feature extraction functions that summarize the recent past using short convolutions or quickly vanishing leaky averages. For instance, the language experiments of Section 7 use feature extractors of the following form: [^2] The leaking average in expression (6) is far too simple to effectively encode long range dependencies as demonstrated in (Voelker et al., 2019; Peng et al., 2023; Gu & Dao, 2023).
[…]
$v_{T}$ $=\mathrm{Norm}\big({\color[rgb]{.75,0,.25}\bar{v}_{T}}\big)$ $\text{with}\quad\underbrace{{\color[rgb]{.75,0,.25}\bar{v}_{T}}={\color[rgb]{0,0,1}\tilde{v}_{T}}+\lambda_{\psi}{\color[rgb]{0,0,1}\tilde{v}_{T+1}}\quad\quad{\color[rgb]{0,0,1}\tilde{v}_{T}}=W_{\psi}\,x_{T}}_{\text{convolution over t=T and T+1}}$
Since this expression produces keys with unit norm ($\mathrm{Norm}(x)=x/\|x\|$), the effective kernel bandwidth is determined by the trainable parameter $\beta$ in equation (3).
@@
Consider now a deep network whose architecture includes layers of associative memory units. When the associative memories are implemented with differentiable kernel smoothing mechanisms, training such a deep network is simply a matter of unrolling the network in time and back-propagating the gradients, in ways that users of modern deep learning software will find very familiar. Unsurprisingly, unrolling equation (3) along an input sequence $(x_{1}\dots x_{D})$ of duration $D$ yields an expression that very much resembles masked self-attention (Vaswani et al., 2017).
$\forall\,T\in\{1\dots D\}\qquad y_{T}=\sum_{i=1}^{T-1}\penalty 10000\ \frac{e^{\beta\,k_{T}^{\!\top}k_{i}}}{\sum_{j=1}^{T-1}e^{\beta\,k_{T}^{\!\top}k_{j}}}\penalty 10000\ v_{i}\penalty 10000\ ,$ (7)
Implementing associative memories with kernel smoothing therefore provides a particularly direct illustration of the connection between self-attention and associative memories (e.g., (Ramsauer et al., 2020)). However, Memory Mosaics differ because the value extraction function is allowed to peek into the near future of the input time series $(x_{t})$. This slight change has important consequences
  • Each memory unit operates as a little predictor whose outputs $y_{T}$ can be interpreted as a conditional expectation (1) that estimates features of the near future ($v_{T}$) of the input time series on the basis of its past observations ($k_{T}$). The parameters of the value extraction function ($\psi$) specify what is being predicted and the parameters of the key extraction function ($\varphi$) specify how it is predicted.
  • Equation (7) must therefore account for the number of future time steps needed to compute $v_{T}$. In our experiments, for example, $v_{T}$ can look one step ahead in the future. This amounts to having a more aggressive attention mask. Therefore the main diagonal must be excluded from the attention mask, justifying the T1{T{-}1} upper bound in the sum. [^3] One could of course use a more aggressive masking to allow $v_{T}$ peeking several time steps in the future.
  • Because each memory unit acts as a predictor, a single layer of memory units is sufficient to address the induction head problem of Bietti et al. (2024). In contrast, a decoding transformer needs at least two self-attention layers for the same task.
  • Equation (7) makes no provision for position encoding and no distinction between query and key vectors. In other words, we are betting that these transformers complications are no longer needed because our associative memory units do not need them to implement induction heads.
4 Predictive Disentanglement
Training and meta-learning
The training process determines which future bit of information is predicted by each associative memory unit (through the parameters that control the computation of the values $v_{T}$) and which kernels are used to perform the predictions (through the parameters of that control the computation of the keys $k_{T}$). In contrast, the relation between keys and predicted values is determined for each input sequence at inference time through the memorization of key/values pairs specific to each sequence. The training procedure should therefore be seen as a meta-learning process, distinct from the memory-based learning that occurs at inference time when new key/value pairs are added into the memories.
Predictive disentanglement
This meta-learning interpretation reveals a remarkable phenomenon that we call predictive disentanglement : the gradient training algorithm splits the overall prediction task (e.g., predicting the next token in a natural language sentence) into disentangled prediction sub-tasks assigned to each memory unit.
Consider a training set composed of long enough sequences $(x_{1},\dots x_{D})$ extracted from underlying time series governed by possibly different stationary processes. The goal of our network is to predict each $x_{T+1}$ using the previous observations $x_{1}\dots x_{T}$. Unrolling the network in time along each sequence $(x_{1}\dots x_{D})$ and collecting the prediction losses measured at each position $t$ can be summarized by a curve that shows the prediction cost (or loss) at each time step $1\dots D$, as illustrated in Figure 2. We can expect that the prediction cost observed at position $T$ becomes smaller when $T$ increases because more information $(x_{1}\dots x_{T})$ is available to predict each $x_{T+1}$.
@@ 4 Predictive Disentanglement
The training process minimizes the total prediction cost, that is the area under the curve in Figure 2 viewed as a collection of vertical slices. We can also view this area as a collection of horizontal slices, each representing the context length required to drive the prediction cost below a certain threshold. Therefore the training process can also be viewed as minimizing the context length needed to produce good enough predictions.
Because the associative memory retrieval function (2) is known to converge to stationary conditional expectations $\mathbb{E}(V|K)$, each memory unit is driven to produce a good conditional expectation estimate as soon as possible. This can be achieved in two ways:
@@
  • Let us first assume that each memory unit has a frozen value extraction function $\psi$. The training procedure can still make each memory unit statistically more efficient by tuning the parameters of the key extraction function $\varphi$, that is, by learning how to compare the current prediction context $(x_{T},x_{T-1},x_{T-2}\dots)$ with past prediction contexts $(x_{t},x_{t-1},x_{t-2}\dots)$ for t<T{t<T}.
@@
    Learning a similarity metric (a kernel) is a well known way to make non-parametric estimators more efficient (e.g., Bach et al., 2004). For instance, the training procedure can construct keys that summarize the relevant contextual information, discarding noise factors that could increase the distance between keys associated with similar values. It can also adjust the effective kernel bandwidth, for instance, using parameter $\beta$ in equation (7).
@@
  • When multiple memory units are available, the training procedure can also distribute the overall prediction task among the available memory units. As long as the memory units outputs can still be combined to address the overall task, the training algorithm can optimize the parameters of the value extraction functions $\psi$ to produce values $v_{T}$ that more efficiently modeled by their respective memory units.
    Because each memory unit operates independently at inference time, this works best when the overall prediction task is disentangled into smaller prediction sub-tasks that can be modeled independently and efficiently. More precisely, the sub-tasks must be chosen so that each memory can carry out its assigned modeling task at inference time without having to account for the combined impact of the operation of all memory units. Their outputs can then be recombined to provide predictions for inputs that are globally very different from the training inputs, but whose disentangled components are individually predictable, as illustrated in Section 5.
@@
Disentanglement has long been recognized as desirable (Bengio, 2013) but has been hard to pinpoint (Comon, 1994; Roth et al., 2022; Thomas et al., 2018). Predictive disentanglement is closely related to the meta-transfer objective of Bengio et al. (2019) but arises as a side effect of a specific predictive architecture trained with the usual gradient procedure. Although predictive disentanglement is easier to understand in the case of a network of associative memory units, we conjecture that something similar also occurs in standard transformers.
@@ 5 Tracking three moons
We give an illustrative example of predictive disentanglement: three moons orbit a remote planet. Although the local astronomers are very far from understanding celestial mechanics, [^4] We do not seek to discuss subtleties such as elliptical orbits or multi-body problems. Our primitive astronomers are best compared to the ancient sky watchers whose efforts eventually gave the Ptolemaic model. they nevertheless observe periodic motions and debate how to predict future moon positions. A first astronomer proposes to compile a single table containing the daily positions of all three moons, arguing that if the current set of moon positions matches a previous observation, the future moon positions will match the following observations. A second astronomer suggests instead to make three tables, one for each moon, arguing that the future positions of each moon can be independently predicted by matching its current position with a previously observed one.
@@
$N_{h}$ $=1\text{\penalty 10000\ or\penalty 10000\ }3$
@@
$y_{t}^{(h)}$ $=\frac{1}{Z_{T}}\,\sum_{t<T}e^{\beta\,k_{T}^{(h)}\bm{\cdot}k_{t}^{(h)}}v_{t}^{(h)}$
@@
[Uncaptioned image] [Uncaptioned image]
Figure 4: Single head network prediction error versus context length. The prediction error shows a sharp transition after $\mathrm{lcm}(p_{1},p_{2},p_{3})$ observations (red vertical line), when the network switches from predicting the future moon position by repeating the last observation to predicting by find a matching memorized configuration.
Figure 5: Three-heads network prediction error versus context length. The prediction error improves whenever the context length reaches the period of a new moon (black vertical lines), yielding accurate predictions after the last one, well before having seen the full set of moon configurations (red vertical line).
@@
For our purposes, each observation $x_{t}$ consists of three complex numbers $e^{i\theta_{k}}$ that encode the angular positions $\theta_{k}$ of the three moons inside their respective orbital plane. We consider two single layer models (Figure 3) with either $N_{h}=1$ or $N_{h}=3$ memory units whose added dimensions match the input dimension. The trainable parameters of the linear key and value extraction are collected in two $3\times 3$} complex matrices $W_{\varphi}$ and $W_{\psi}$. The memory unit follow equation (3) with a fixed parameter $\beta=50$. A third $3\times 3$} complex matrix $W_{z}$ combines the memory unit predictions into an output $z_{T}$ that hopefully predicts $x_{T+1}$. Both networks share an interesting analytic solution: setting all three matrices $W_{\varphi}$, $W_{\psi}$, and $W_{z}$ to the identity yields optimal predictions once the associative memories have seen enough samples.
@@
Figure 5 and 5 show the prediction errors of both networks as a function of the context length, that is, the number of observations stored into the memories. More precisely, for each sequence $(x_{t})$ and each time index $T$, we compute the average absolute deviation between the next 25 true moon positions $x_{T+1}\dots x_{T+25}$ and the next 25 auto-regressive predictions (in which the successive predictions are looped back into the network input.) The plots show curves averaged over 512 sequences sharing the same set of moon periods taken from either the training or validation set.
@@
  • For the three-heads network (Figure 5), the prediction error curve drops after seeing exactly $p_{1}$, $p_{2}$, and $p_{3}$ observations (black lines), that is whenever the orbit of an additional moon has been memorized. The learned weight matrices are shown Figure 10 in the Appendix. Observe how the network produces accurate predictions after a time equal to the period $\max(p_{1},p_{2},p_{3})$ of the slowest moon (last black line), long before the combined period $\mathrm{lcm}(p_{1},p_{2},p_{3})$ (red line) of the moon system. In this interval, accurate predictions are returned for moon configurations that can be very different from the previously observed ones. Instead the network combines individual moon predictions, each well supported by the past observations.
Predictive disentanglement and compositional learning in language models
Consider a chat-bot assisted creative writing scenario in which the human uses dialogue to repeatedly introduce new ideas into an evolving story that the chat-bot reprints at each step. The user can drive such a story arbitrarily far from the training data and into the distant tail of its distribution. Although no training example resembles the story, the chat-bot keeps producing syntactically correct language and coherent stories because it has learned some of the mathematical structures of language (Harris, 1968) and can recombine pieces of information coming from either the context or the training data. This phenomenon is fundamentally similar to that illustrated in Figure 5, where moon configurations unlike any previously seen configurations are accurately predicted because the network has learned how to combine individual moon predictions. This similarity casts a useful light on the otherwise mysterious compositional learning abilities of transformer-like models.